OpenStudy (anonymous):

The perimeter of a rectangle is 90 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 8: 7. 30 ft by 15 ft 26 ft by 19 ft 24 ft by 21 ft 20 ft by 25 ft

OpenStudy (anonymous):

Perimeter = 2(l+w) l /w = 8/7 => l = 8/7w 90 = 2(w+8/7w) Solve for w then for l

OpenStudy (anonymous):

???

OpenStudy (catlover5925):

ok so I solved for W and I got 14.5 it is closest to A so that would be my answer

OpenStudy (anonymous):

And W means width right?

OpenStudy (catlover5925):

right

OpenStudy (anonymous):

Ok

OpenStudy (catlover5925):

did I help at all?

OpenStudy (anonymous):

How did u find the width

OpenStudy (anonymous):

The perimeter divided by what?

OpenStudy (anonymous):

90 = 2(w +8/7w) 45 = 15/7w 45*7 /15 = w 21 = w l = 8/7 w l = 24

OpenStudy (anonymous):

How did you get 15/7w

OpenStudy (anonymous):

w + 8/7w = (7w +8w)/7

OpenStudy (anonymous):

SO basically 7+8 divided by 7??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2.1428... is what i got ??

OpenStudy (anonymous):

w is still unidentified here. you have 45 = (8+7)/7 * w

OpenStudy (anonymous):

WOAH...where did u get 45?!?!

OpenStudy (anonymous):

Im SOO sorry...i gotta go, Bye :(

OpenStudy (anonymous):

Perimeter of a rectangle is P = 2(l+w) l/w = 8/7 (given in question) so l = 8/7 w Perimeter = 90 (given ) 90 = 2(w +8/7w) 90/2 = (7+8)/7 * w 45 = 15/7 *w